OpenStudy (anonymous):

y=-x+5 into standard form?

OpenStudy (anonymous):

the standard for is of the form ax+by+c=0 convert it into that form and you will get the answer

OpenStudy (anonymous):

Did you get thw answer @AlleyPony

OpenStudy (anonymous):

i had gotten -y-x=5, but wasnt sure if it was that or y+x=5

OpenStudy (anonymous):

Both the answer are not correct... Take allt eh terms on L.H.S of the equation and you will get the answer @AlleyPony

OpenStudy (anonymous):

Your answer should be x+y-5=0

OpenStudy (anonymous):

oh, okay. thank you. I didn't realize it had to be zero!

OpenStudy (anonymous):

I thought it was Ax+By=C?